OpenStudy (anonymous):

find derivative: f(u) = arctan(u/(1+u)

sam (.sam.):

1/(2u^(2)+2u+1)

OpenStudy (anonymous):

oh is it because you need to foil them out?

sam (.sam.):

arctan u/(1+u) Remove the factor of 1 from the numerator. 1/(v^(2)+1) Use the quotient rule to find the derivative of u/(1+u). The quotient rule states that ((f)/(g))'=(f'g-fg')/(g^2)' d/dx [u]((1+u))-(u)((d)/(dx) [(1+u)]))/(((1+u))^(2)) Simplify the derivative. d/dx u/(1+u)=1/(1+u)-u/(1+u)^(2) Remove the factor of 1 from the numerator. (1)/(1+u)-(u)/((1+u)^(2)) Simplify the derivative. d/du arctan u/(1+u)=1/(2u^(2)+2u+1) Remove the factor of 1 from the numerator. 1/(2u^(2)+2u+1)

OpenStudy (anonymous):

ok i think i get it but then how about this one? f(x) = cos(arcsin(x+1))

sam (.sam.):

\[\cos(\arcsin(x+1))\] Add 0 to 1 to get 1 \[\frac{d}{dx}x+1=1\] Simplify the derivative \[\frac{d}{dx}\arcsin(x+1)=\frac{1}{\sqrt{-x(2+x)}}\] Simplify the derivative \[\frac{d}{dx}\cos(\arcsin(x+1))=-\frac{\sin((\arcsin(x+1))}{\sqrt{-x(2+x)}}\] The derivative of cos(arcsin(x+1) is \[-\frac{\sin((\arcsin(x+1))}{\sqrt{-x(2+x)}}\]

OpenStudy (anonymous):

thanks! :)